The width of a rectabgle is 1 meter more than half the length of the rectangle. If the perimeter of the rectangle is 40 meters, find the width and the length. I know A=w*l and P=2l+2w, however I can not figure out how to set this problem up.
Thank you,
Feona

Answer by cleomenius(959)

length = x
width = 1/2x + 1
2(x) + 2(1/2x + 1) = 40
This is the part I think you needed help with; setting up the equation.
=======================================================================
2x + x + 2 = 40
3x = 38
x = 12.7 meters = length
7.35 meters = width
2(12.7) + 2(6.35 + 1 ) = 40 meters.
25.4 + 12.7 + 2 = 40.1 (round off error)
Cleomenius.